# Machine learning applications in archaeological practices: a review

### v2

@@ 4.2.2 Archaeological predictive models (APMs)
Already present since the development of settlement patterns analysis (Willey, 1953), the tradition of APMs rose alongside the development of modelling and prediction theory in archaeology (Judge and Lynne, 1988). It is therefore surprising to see little interest in applying machine learning methods to this approach in recent years even down to 2022 [Fig. 5A]. Critics can point to the relative reluctance of researchers towards APMs [Kvamme 2006, p.7, Table 1.1, Lock and Harris 2006, pp. 42-45]. One main criticism about APMs is that “… predictive modelling as it is presently practiced is fundamentally about environmental determinism.(Kohler, 1988, p.19). This statement is particularly true when we consider the strong influence of environmental and ecological studies in the field of landscape archaeological (cf. 3.3.4). However, even if this criticism has neither recent origins, nor is it confined to the past (Wheatley, 2004; Arponen et al., 2019; Kristiansen, 2019), responses to it exist. Coombes and Barber emphasise the interest of models despite their incorrectness, stating that ”A simple model cannot hope to replicate all the complexities of environment–culture relationships across a civilization, but one basic approach that can provide valuable insights is to treat human populations in ecological terms, with their ranges shifting in response to changing conditions(Coombes and Barber, 2005, p. 305). Furthermore, new applications of APMs dedicated to cultural heritage preservation against looting (El-Hajj, 2021) can prove to be a promising solution to further aid in the protection of cultural heritage alongside more intensive public funding.
@@
The frequent use of the MaxEnt model for APMs, initially designed by ecologists (Phillips et al., 2006), shows the influence of environmental niche concepts in archaeology (Demjan et al., 2022; Vernon et al., 2022; Yaworsky et al., 2020). MaxEnt is well suited to examine questions of dispersion and provides more robust and deterministic archaeological predictions due to its incorporation of absence of data into its modelling: “As a result, MaxEnt is more suitable for archaeological data than the other predictive modelling approaches(Yaworsky et al., 2020, p. 14).
@@
An essential consideration in APMs is the number of covariates, predictors, or variables required for the building of the models. In Yaworsky et al. [2020], 55 predictors were used, Hansen and Nebel [2020] employed 26 predictors, Friggens et al. [2021] used 25 predictors, Castiello and Tonini [2021] used 13 predictors, and Benner et al. [2019] relied on eight predictors. However, if data transformation techniques such as PCA are applied, which seek to reduce the number of covariates whilst maximising the impact of those that remain, a smaller number of predictors can be used for the final model instead (Hansen and Nebel, 2020; Yaworsky et al., 2020; El-Hajj, 2021).
@@ 4.2.3 Automatic structures detection
Over the past decade, there has been an explosion in machine learning methods applied to automatic structure detection tasks [Fig. 5A]. Though earlier applications were mostly based on satellite STRM or ASTER images (Menze et al., 2006; Menze and Ur, 2012), the recent explosion of light detection and ranging (LiDAR) and airborne laser scanning (ALS) data (Wurzer et al., 2015; Gillings et al., 2020), has led to an increase in interest for this field. While earlier studies employed random forest methods (Menze et al., 2006; Guyot et al., 2018; Stott et al., 2019), recent papers mainly applied ANNs, which aligns with the predominance of image data and neural network architectures that excel at image classification tasks. The recent multi-scale region-based convolutional neural network (MR-CNN) model, adapted from convolutional neural networks (CNNs), has been eagerly applied by archaeologists, as its raison d’être is well aligned with the task of detecting archaeological structures from image data (Bundzel et al., 2020; Bonhage et al., 2021; Davis et al., 2021; Guyot et al., 2021; Altaweel et al., 2022; Banasiak et al., 2022; Fisher et al., 2022). The MR-CNN architecture also allows for image segmentation by masking the detected region of interest. This architecture substantially helps researchers recognise archaeological structures, as well as verify the model’s prediction. Furthermore, the emergence of transfer learning has amplified the use of CNNs in archaeology (Gallwey et al., 2019; Soroush et al., 2020; Herrault et al., 2021). By using a neural network model trained for a different, but similar, type of image data (e.g. to detect bicycles in an image), and training it on the actual desired data (e.g. detect motorcycles) but with a much smaller dataset, compared to the dataset needed without transfer learning, transfer learning can be helpful where only small datasets are available, reducing thus also the time and cost of data acquisition and preparation. In our dataset of articles, the use of transfer learning – counted at 41% (n=16) of all study cases reviewed on automatic structure detection - were generally based on models pre-trained with image datasets such as ImageNet (Russakovsky et al., 2015) or COCO (Lin et al., 2014), though in one unique case authors also relied on specialised datasets (Silburt et al., 2019).
@@
To evaluate the success of the machine learning archaeological structure detection model model, authors have access to a wide range of metrics. F1-score is the most widely used with 20 accounted uses [e.g. Character et al., 2021; Altaweel et al., 2022; Banasiak et al., 2022], but prediction accuracy is also common, with 16 application in the study cases on automatic structure detection [e.g. Monna et al. 2020; Davis and Lundin 2021]. While not always directly discussed, precision, average precision, and recall can nevertheless be found in the results in most of the study cases reviewed here (Bonhage et al., 2021; Berganzo-Besga et al., 2021). Intersect over union (IoU) is only present in five study cases (Bundzel et al., 2020; Bordon et al., 2021; Banasiak et al., 2022; Trotter et al., 2022; Yang et al., 2022), despite revealing valuable information on a model accuracy and being particularly well-suited for image segmentation tasks. The newly developed “boundary IoU” (Cheng et al., 2021) was specifically created for evaluating image segmentation models , and could become an important addition to future studies. New metrics have even been developed explicitly for archaeological image segmentation. Fiorucci et al. [2022] introduced centroid-based and pixel-based measures, as they found IoU not ideal for evaluating discrete archaeological objects.
The significant popularity of machine learning applications for automatic structure detection can be further found in its perceived suitability to answer the research question of whether or not a certain feature is an archaeological site or structure?”. Machine learning methods seem to be particularly effective when applied to massive structures such as burial mounds (Guyot et al., 2018; Caspari and Crespo, 2019; Monna et al., 2020; Berganzo-Besga et al., 2021) or large-size archaeological sites (Menze et al., 2006; Menze and Ur, 2012; Stott et al., 2019)).
@@ 4.2.4 Digital heritage
The high number of ANN models for semantic image segmentation adapted to cultural heritage tasks confirms a tendency (Sultana et al., 2020), with ANNs generally obtaining better results compared to decision trees (Boston et al., 2022). Non-standardised datasets and the diversity in the preprocessing stages made for a heterogeneous assemblage of reviewed study cases, making them difficult to adequately compare. Furthermore, the absence of standardised protocols led to the application of older models (Fiorucci et al., 2020, part. 7) that do not take advantage of the advances in computer vision and semantic segmentation [e.g. Pepe et al. 2022]. Limiting the use of colour images or reducing the number of classes has been suggested to improve future models (Grilli and Remondino, 2019).
@@ 4.2.6 Taphonomic classification
All articles found were co-authored by either Manuel Domínguez-Rodrigo or Enrique Baquedano, or other authors that nevertheless contributed with them. No other articles by a workgroup unaffiliated with either author were found in our searches.
Many of the articles applied a large number of machine learning methods as well as methods that we do not consider machine learning [such as PLSDA and MDA; e.g. Domínguez-Rodrigo and Baquedano 2018; Abellán et al. 2022] to compare their performance on the specific task, whilst others [e.g. Byeon et al. 2019; Cifuentes-Alcobendas and Domínguez-Rodrigo 2019] were more restrained. This meant that there was no clear machine learning method or family of methods that was used much more frequently than others, although neural networks were applied in all but one of the articles reviewed (cf. 3.3.8), which is not the case for other methods also commonly used in these articles, such as SVMs or random forests.
@@
However, despite the wide range of methods, many of the articles reviewed were found to contain important methodological issues, both from unclear and insufficient reporting of the methods applied, to important considerations of the data used that were seemingly eschewed.
Some of these methodological issues have already been highlighted previously [McPherron et al., 2022; but see the reply by Abellán et al. 2022; Moclán and Domínguez-Rodrigo 2023], especially dealing with the unclear methodology applied for the process of “bootstrapping” described in one publication also part of our reviewed dataset [Domínguez-Rodrigo 2018; though note that this is also present in other publications in our dataset: Domínguez-Rodrigo and Baquedano 2018; Aramendi et al. 2019; Moclán et al. 2019, 2020]. The lack of specificity on how this process was undertaken is serious, since bootstrapping to create a larger overall dataset and then splitting data for validation and testing [as seems to be very explicitly stated in Domínguez-Rodrigo and Baquedano 2018; Domínguez-Rodrigo 2018; Aramendi et al. 2019] constitutes a methodological failure, as all the machine learning models will have the same data for both training and testing datasets, rendering any validation results unusable in the same way providing an exam to a student together with its answer sheet would. Although we do not wish to reiterate the discussion on the use or misuse of the term “bootstrapping” from Domínguez-Rodrigo [2018], it is important to highlight how unclear reporting of methodology or code unavailability can lead other researchers to consider the methods applied unsound. For another related example, both Domínguez-Rodrigo [2018] and Domínguez-Rodrigo and Baquedano [2018] report the use of bootstrapping, but then also describe the use of k-fold cross validation performed seemingly alongside it, both of which are part of the caret R package (Kuhn, 2019), but cannot be used together, as seems to be implied. The trainControl command of the caret package is presumably the one used in both articles, even though no code was made available, and no specific package version was specified [except in an appendix of a citation Kuhn 2019, p.556] to confirm any possible changes in the command’s behaviour in the past that would have made it possible. A more generous reading is that this perhaps only applies to one (C5.0 decision trees) of the six algorithms used, for which k-fold cross validation is specifically mentioned, as can be perhaps surmised from Aramendi et al. [2019], who only mention k-folds cross validation specifically for C5.0. Even so, this raises the question of whether k-folds was performed in addition to holdout testing for all methods [as seems to be the case in Domínguez-Rodrigo and Baquedano 2018; Domínguez-Rodrigo 2018], or if the C5.0 implementation alone did not use holdout testing, or was the only one to use both [either option possible in Aramendi et al. 2019]. This also exemplifies the importance of carefully reporting the software versions of the programs used, if not making code available outright.
@@
Other issues may require more technical knowledge to identify, such as those in Moclán et al. [2019], another taphonomic classification study suffering from the main bootstrapping issue discussed above, though only partially, as they evaluated all the models used with both a bootstrapped and non-bootstrapped sample. This study sought to predict the agent that caused a specific bone fracture plane in bone fragments (in which many planes can be present) made experimentally; in this case, between anthropic direct percussion, spotted hyena (C. crocuta) carnivory and wolf (C. lupus) carnivory (Moclán et al., 2019). However, the bone fracture plane entries dataset of fracture planes reported in the article’s Electronic Supplementary Material 1 (ESM 1) contain information on the bone fragment they are from, rather than only about the fracture plane itself (Moclán et al., 2019, app. ESM 1). Thus, if the dataset were randomly split, some data points from the same bone fragment used for model validation may have contained identical bone fracture metadata to the training input, allowing the model to use that metadata to predict the agent of the fracture plane, since all bone fracture planes from the same bone fragment were caused by the same agent (cf. Varoquaux and Cheplygina 2022). In some cases a single bone fragment had six fracture plane data points defined (and thus six data points available to be shuffled into both testing and training), all containing the same bone fragment metadata for the variables of epiphysis presence, length, and number of planes (Moclán et al., 2019, app. ESM 1). Though no code was made available for this publication to examine how these metadata variables were used during model training and validation, the authors do state that they were indeed used (Moclán et al., 2019, p.4668). In addition, the number of entries in ESM 1 do not match the number of fragments reported for any of the three classes (Moclán et al., 2019, app. ESM 1). Even when counting only unique bone fragments instead of planes (using ESM 1), the numbers, while closer to the numbers reported in the Results section, still do not match for any class (Moclán et al., 2019). Furthermore, for the Hyena class, it is also clear some fragments did not have all their fracture planes added to the ESM 1 dataset [e.g. there are an odd number of entries that state the planes per bone fragment is 2, thus implying there is one plane that was not included; Moclán et al. 2019].
The issues described above for Moclán et al. [2019] may even be present in another article in this category (Moclán et al., 2020), though the latter eschewed the use of bootstrapping altogether, as well as any supplementary material. Despite not publishing the dataset or code, the authors stated that the article was a continuation of Moclán et al. [2019], citing it continuously as their methodological starting point, and likely using the same model trained in Moclán et al. [2019], especially as the variables described were identical. Moclán et al. [2020] focused on classifying bone fracture agency in archaeological – as opposed to experimental – material [contrary to Moclán et al. 2019], though without correcting any of the issues described above, despite providing strong conclusions on the agent that caused the archaeological bone fracture based on the results of a model with considerable unaddressed issues. The use of machine learning methods to draw any conclusion from its own predictions of archaeological data must be very carefully done, the limitations of the methods must be clearly and transparently examined, and the strength of the conclusions must be scrupulously tempered. Anything else should not be considered good scientific practice.
@@
Despite the earlier discussion, however, not all study cases reviewed had such major methodological issues. The methods used in other articles in this category such as Byeon et al. [2019] and Cifuentes-Alcobendas and Domínguez-Rodrigo [2019] are promising and well-grounded (with only a few caveats, such as the latter’s omission on whether BSMs from the same bone were split into both training and testing, which could have unduly increased validation accuracy) and should definitely be the basis of further studies on the subject.

### v3

@@ 4.2.2 Archaeological predictive models (APMs)
Already present since the development of settlement patterns analysis (Willey, 1953), the tradition of APMs rose alongside the development of modelling and prediction theory in archaeology (Judge and Sebastian, 1988). It is therefore surprising to see little interest in applying machine learning methods to this approach in recent years even down to 2022 [Fig. 5A]. APMs focus on predicting unknown or undiscovered sites so that they can be discovered and documented based on previous field observations and variable predictors, for example climate, topography and anthropic features Brandt et al. (1992); Kvamme (2006). Critics can point to the relative reluctance of researchers towards APMs [Kvamme 2006, p.7, Table 1.1, Lock and Harris 2006, pp. 42-45]. One main criticism about APMs is that “… predictive modelling as it is presently practiced is fundamentally about environmental determinism.(Kohler, 1988, p.19). This statement is particularly true when we consider the strong influence of environmental and ecological studies in the field of landscape archaeology (cf. 3.3.4). However, even if this criticism has neither recent origins, nor is it confined to the past (Wheatley, 2004; Arponen et al., 2019; Kristiansen, 2019), responses to it exist. Coombes and Barber emphasise the interest of models despite their incorrectness, stating that ”A simple model cannot hope to replicate all the complexities of environment–culture relationships across a civilization, but one basic approach that can provide valuable insights is to treat human populations in ecological terms, with their ranges shifting in response to changing conditions(Coombes and Barber, 2005, p. 305). Furthermore, new applications of APMs dedicated to cultural heritage preservation against looting (El-Hajj, 2021) can prove to be a promising solution to further aid in the protection of cultural heritage alongside more intensive public funding.
@@
The frequent use of the MaxEnt model for APMs, initially designed by ecologists (Phillips et al., 2006; Sillero et al., 2021), shows the influence of environmental niche concepts in archaeology (Demján et al., 2022b; Vernon et al., 2022; Lundström et al., 2024; Yaworsky et al., 2024a, b). MaxEnt is well suited to examine questions of dispersion and provides more robust and deterministic archaeological predictions due to it being a presence-only model, using background points rather than true absence or pseudo-absence data (used by other methods, including machine learning) into its modelling: “As a result, MaxEnt is more suitable for archaeological data than the other predictive modelling approaches(Yaworsky et al., 2020, p. 14). Furthermore, when compared to other models such as random forest, which can be prone to overfitting, MaxEnt appears as the best option for species dispersion problems (Valavi et al., 2022).
@@
An essential consideration in APMs is the number of covariates, predictors, or variables required for the building of the models. In Yaworsky et al. [2020], 55 predictors were used, Hansen and Nebel [2020] employed 26 predictors, Friggens et al. [2021] used 25 predictors, Castiello and Tonini [2021] used 13 predictors, and Benner et al. [2019] relied on eight predictors. However, if data transformation techniques such as PCA are applied, which seek to reduce the number of covariates whilst maximising the impact of those that remain, a smaller number of predictors can be used for the final model instead (Hansen and Nebel, 2020; Yaworsky et al., 2020; El-Hajj, 2021). Research on species distribution models assess the maximum number of predictors at $k=\frac{(n-50)}{8}$ (Field et al., 2012), $k$ being the number of predictors and n the number of occurrences of the species.
@@ 4.2.3 Automatic structures detection
Over the past decade, there has been an explosion in machine learning methods applied to automatic structure detection tasks [Fig. 5A]. Though earlier applications were mostly based on satellite STRM or ASTER images (Menze et al., 2006; Menze and Ur, 2012), the recent explosion of light detection and ranging (LiDAR) and airborne laser scanning (ALS) data (Wurzer et al., 2015; Gillings et al., 2020), has led to an increase in interest for this field. While earlier studies employed random forest methods (Menze et al., 2006; Guyot et al., 2018; Stott et al., 2019), recent papers mainly applied ANNs, which aligns with the predominance of image data and neural network architectures that excel at image classification tasks. The recent multi-scale region-based convolutional neural network (MR-CNN) model, adapted from convolutional neural networks (CNNs), has been eagerly applied by archaeologists, as its raison d’être is well aligned with the task of detecting archaeological structures from image data (Bundzel et al., 2020; Bonhage et al., 2021; Davis et al., 2021; Guyot et al., 2021; Altaweel et al., 2022; Banasiak et al., 2022; Fisher et al., 2022). MR-CNNs segment images via masked regions, aiding structure identification and validation. Furthermore, the emergence of transfer learning has amplified the use of CNNs in archaeology (Gallwey et al., 2019; Soroush et al., 2020; Herrault et al., 2021). By using a neural network model trained for a different, but similar, type of image data (e.g. to detect bicycles in an image), and training it on the actual desired data (e.g. detect motorcycles) but with a much smaller dataset, compared to the dataset needed without transfer learning, transfer learning can be helpful where only small datasets are available, reducing thus also the time and cost of data acquisition and preparation. In our dataset of articles, the use of transfer learning – counted at 41% (n = 16) of all study cases reviewed on automatic structure detection - were generally based on models pre-trained with image datasets such as ImageNet (Russakovsky et al., 2015) or COCO (Lin et al., 2014), though in one unique case authors also relied on specialised datasets (Silburt et al., 2019).
@@
To evaluate the success of machine learning archaeological structure detection models, authors have access to a wide range of metrics. F1-score is the most widely used with 20 accounted uses [e.g. Character et al., 2021; Altaweel et al., 2022; Banasiak et al., 2022], but prediction accuracy is also common, with 16 applications in the study cases on automatic structure detection [e.g. Monna et al. 2020; Davis and Lundin 2021]. While not always directly discussed, precision, average precision, and recall can nevertheless be found in the results in most of the study cases reviewed here (Bonhage et al., 2021; Berganzo-Besga et al., 2021). Intersect over union (IoU) is only present in five study cases (Bundzel et al., 2020; Bordon et al., 2021; Banasiak et al., 2022; Trotter et al., 2022; Yang et al., 2022), despite revealing valuable information on a model accuracy and being particularly well-suited for image segmentation tasks. The newly developed “boundary IoU” (Cheng et al., 2021) was specifically created for evaluating image segmentation models, and could become an important addition to future studies. New metrics have even been developed explicitly for archaeological image segmentation. Fiorucci et al. [2022] introduced centroid-based and pixel-based measures, as they found IoU not ideal for evaluating discrete archaeological objects.
The significant popularity of machine learning applications for automatic structure detection can be further found in its perceived suitability to answer the research question of whether a certain feature is an archaeological site or structure?. Machine learning methods seem to be particularly effective when applied to massive structures such as burial mounds (Guyot et al., 2018; Caspari and Crespo, 2019; Monna et al., 2020; Berganzo-Besga et al., 2021) or large-size archaeological sites (Menze et al., 2006; Menze and Ur, 2012; Stott et al., 2019)).
To summarise, following the explosion of big data such as the use of archaeological remote sensing imagery (especially LiDAR), combined with the development of powerful deep learning models, new possibilities for the identification of settlements and archaeological structures have appeared. Leaning massively on transfer learning and pre-trained models, the aforementioned studies applying these newly developed models provide encouraging results to archaeologists. However, the diversity of applications, metrics, pre-processing, and the very nature of the neural network models all lead to a heterogeneity of methodologies available for this task.
@@ 4.2.4 Digital heritage
The high number of ANN models for semantic image segmentation adapted to cultural heritage tasks confirms a tendency (Sultana et al., 2020), with ANNs generally obtaining better results compared to decision trees (Boston et al., 2022). Non-standardised datasets and the diversity in the preprocessing stages made for a heterogeneous assemblage of reviewed study cases, making them difficult to adequately compare. The lack of standard protocols (Fiorucci et al., 2020, part. 7) has led to the continued use of older models [e.g. Pepe et al. 2022]. Limiting the use of colour images or reducing the number of classes has been suggested to improve future models (Grilli and Remondino, 2019).
Two different approaches are currently used in machine learning techniques for digital heritage problems: one emphasises a more progressive and integrative workflow with, successive detailed pre-processing steps, while the other focuses on semantic segmentation. Although deep learning models, in particular ANNs, appear to be the most used models, results are still hardly trustworthy in most cases, due to the lack of comparative datasets and standardised metrics.
@@ 4.2.5 Text analysis
Despite its essential role in archaeological research, machine learning applications for text analysis are poorly represented in our corpus. Almost only relying on ANN models in our corpus, these applications face strong methodological and technical issues during the pre-processing of the textual data. However, the recent development of large language models, absent in our corpus, is likely to drastically change the automated analysis of archaeological textual data in the future.
@@ 4.2.6 Taphonomic classification
Despite the wide range of methods applied, many of the articles reviewed were found to contain important methodological issues, both from unclear and insufficient reporting of the methods applied, to important considerations of the data used that were seemingly eschewed.
Some of these concerns have already been highlighted [McPherron et al., 2022; but see the reply by Abellán et al. 2022, including its Supplementary Texts 1 and 2; Moclán and Domínguez-Rodrigo 2023; see also Courtenay et al. 2024], specifically the unclear methodology for the process of "bootstrapping” described in one publication in our reviewed dataset [Domínguez-Rodrigo 2018; but also present in our dataset in Domínguez-Rodrigo and Baquedano 2018; Aramendi et al. 2019; Moclán et al. 2019, 2020; and Courtenay et al. 2019]. The unclear methodology reported in Domínguez-Rodrigo [2018] could be interpreted as stating that bootstrapping was applied to create a larger overall dataset which was then split for training and testing (as seems to also be explicitly stated in Domínguez-Rodrigo and Baquedano 2018; but less so in Aramendi et al. 2019, or Courtenay et al. 2019]. This constitutes a methodological failure, as the models will share the same data for across both training and testing datasets, rendering any results unusable in the same way it would be to have an answer sheet when taking an exam. Although we do not wish to reiterate the “bootstrapping” discussion, it is still important to highlight how imprecise reporting of methodology and code unavailability can cast doubt of the validity of the methods applied.
When the data and code is made available, such as in Abellán et al. [2022], concerns are easier to both discard and confirm. For instance, the Ensemble.ipynb file provided in the Supplementary Data (henceforth SD) 2 contains an off-by-one error, which has the effect of excluding the last variable when training the models used. If this is indeed the code used for the main article (it is at least for the Supplementary Text 4), it would have considerable implications for the reported results and the accuracy of the reported methods. Moreover, the R file croc paper 2022 code.r contains additional problems for the “correct” bootstrapping section (module 4 in the file), and we could not run the code for this module without errors (with R 4.3.2, and caret 7.0-1).
On the other hand, the code suggests that the definition of "bootstrapping” used by the authors is to resample input data with replacement to obtain a larger synthetic input dataset and then performing k-folds when training the model (perhaps explaining the reported use of both bootstrapping and k-folds in Domínguez-Rodrigo 2018 and Domínguez-Rodrigo and Baquedano 2018). This is contrary to the ideal implementation of bootstrapping, where the resampling is done by each classifier in the ensemble learning model, thus creating multiple bootstrapped datasets (one per classifier), as well as providing an out-of-bag error to verify the accuracy of each classifier when aggregating the models (Breiman, 1996). The code also implies that their "proper bootstrapping” process for creating a larger dataset may also be applied to the testing data after splitting as well, but this would severely skew model accuracy metrics, since the same data could be correctly or incorrectly predicted multiple times.
@@
Other issues may require even more technical knowledge to identify, such as those in Moclán et al. [2019], who evaluated their models with both a bootstrapped and non-bootstrapped experimental sample, seeking to predict the agent (anthropic direct percussion, spotted hyena (C. crocuta) carnivory and wolf (C. lupus) carnivory) of bone fracture planes in bone fragments (Moclán et al. 2019). However, the dataset provided in the Electronic supplementary material 1 (ESM 1) also contains information on the bone fragment the fracture planes come from (e.g. epiphysis presence, length, and number of planes), rather than only the information of the plane itself (Moclán et al., 2019, app. ESM 1). In some cases a single bone fragment could have six fracture planes in the dataset, so when splitting the dataset, some input data may have come from the same bone as some testing data, and thus have this shared bone metadata, likely leading to overfitting as all bone fracture planes from the same bone fragment were caused by the same agent [see (Varoquaux and Cheplygina, 2022)]. Though no code was made available, the authors do state that these metadata variables were indeed used (Moclán et al., 2019, p. 4668). In addition, the number of entries in ESM 1 do not match the number of fragments reported for any of the three classes. Even when counting only unique bone fragments instead of planes using ESM 1, the numbers still do not match for any class. Furthermore, for the hyena class, it is also clear some fragments did not have all their fracture planes added to the ESM 1 dataset, since there are an odd number of entries that state the planes per bone fragment is 2, thus implying there is at least one plane that was not included (Moclán et al., 2019).
The issues in Moclán et al. [2019] are likely also present in another article [Moclán et al. 2020; and perhaps by Abellán et al. 2022, at least for the tramp_cm_croc_MDR_EB.txt dataset, which contains a "number of grooves" variable). Although the data and code used by Moclán et al. 2020 was not made public, the authors state that the work was a continuation of Moclán et al. [2019], repeatedly citing it as their methodological starting point, and likely using its trained model, especially as the variables described were identical. Moclán et al. [2020] focused on classifying bone fracture agency in archaeological – as opposed to experimental – material, They provide strong conclusions based on the results of a model with considerable unaddressed (and unacknowledged) problems. A similar issue could be raised for Abellán et al. [2022], though they are far more couched in their language. Nevertheless, Fig. 4 in Abellán et al. [2022] states that for some of the archaeological BSMs analysed, the models gave a classification probability of 99% for being crocodile-made, but at the same time also a 99% probability for being a cut mark. Oddly, the probabilities do not add up to 1 except for the last BSM shown (BOU-VP-11/12), the only one classified as a cut mark. Whether this was a transcribing error or a deeper issue is unclear, as detailed results are not available in the SD. The vast class imbalance of input data is another issue, despite the authors suggesting otherwise. Even classifying no crocodile BSMs correctly would provide an F1 score of 0.92 if only 11 out of 146 testing cut marks were misclassified, thus making the lack of confusion matrix or detailed results reporting concerning, even if no strong conclusion was put forth [see Courtenay et al. 2024].
The use of machine learning methods to draw any conclusion from its own predictions of archaeological data must be very carefully done, the limitations of the methods must be clearly and transparently examined, and the strength of the conclusions must be scrupulously tempered. Anything else should not be considered good scientific practice.
@@
However, despite the lengthy discussion above, not all study cases reviewed had methodological shortcomings. The methods used in other articles in this category such as Byeon et al. [2019] and Cifuentes-Alcobendas and Domínguez-Rodrigo [2019] are promising and well-grounded [with only a few caveats, such as the latter’s omission on whether BSMs from the same bone were split into both training and testing, which could have unduly increased validation accuracy; see also Courtenay et al. 2024] and should definitely be the basis of further studies on the subject, including those seeking to independently verify and replicate the results of the numerous BSM agent classification studies, especially as recent attempts have not obtained the same optimistic results, despite employing robust methodology (Courtenay et al., 2024).